OpenStudy (anonymous):

The speed of sound in ocean water is about 1,530 m/s. If it takes 3 seconds for a sound to travel from a ship to the bottom of the ocean and back, how deep is the water?

OpenStudy (wasiqss):

1530*3

OpenStudy (wasiqss):

4590 m

OpenStudy (anonymous):

can you give more detail??

OpenStudy (wasiqss):

distance =speed *time hence= 1 530*3 speed is 1530 , time is 3

OpenStudy (anonymous):

its 1530*1.5= 2295 metres

OpenStudy (anonymous):

D = S*T 2x=S*T 1,530*3 x = (1530*3)/2

OpenStudy (wasiqss):

sorrry did wrong oops

OpenStudy (anonymous):

I need one lol,, and detail??

OpenStudy (wasiqss):

dint read full question

OpenStudy (wasiqss):

nahh its is coming back to in 3 second so two journey is 3 seconds, so one journey in 1.5 seconds

OpenStudy (lilg132):

1530 * 3 = 4590 / 2 = 2295 metres deep

OpenStudy (anonymous):

1530 meters / 1 sec = X meters / 1.5 sec (it's a ratio) then solve for X, and you get 2295 meters.